2 LINEAR EQUATIONS IN 2 UNKNOWNS or SIMULTANEOUS LINEAR EQUATIONS IN 2 UKNOWNS.
twice the larger of two number plus the smaller is -25. The larger minus three times the smaller is 138. find the smaller number

Answer by ankor@dixie-net.com(22740)

twice the larger of two number plus the smaller is -25.
2x + y = -25
The larger minus three times the smaller is 138.
x - 3y = 138
rearrange to
x = 3y + 138
find the smaller number
replace x with (3y+138) in the first equation
2(3y+138) + y = -25
6y + 276 + y = -25
6y + y = -25 - 276
7y = -301
y = -301/7
y = -43 is the smaller number